A clinical trial is conducted comparing a new pain reliever for arthritis to a placebo. Participants are randomly assigned to receive the new treatment or a placebo. The outcome is pain relief within 30 minutes. The data are shown in table 6-8

Pain relief no Pain relief
New medication 44 76
Placebo 21 99

a. Generate a 95% confidence interval for the proportion of patients on the new medication who report pain relief.
b. Generate a 95 % confidence interval for the difference in proportions of patients who report pain relief.

how it calculated with 30 mins ? I'm still not understand.

Answer by stanbon(57377)

A clinical trial is conducted comparing a new pain reliever for arthritis to a placebo. Participants are randomly assigned to receive the new treatment or a placebo. The outcome is pain relief within 30 minutes. The data are shown in table 6-8

Pain relief no Pain relief
New medication 44 76
Placebo 21 99

a. Generate a 95% confidence interval for the proportion of patients on the new medication who report pain relief.
sample proportion: 65/240 = 0.27
standard error = E = 1.96*sqrt[0.27*0.73/240] = 0.056
---
95% CI: 0.27-0.056 < p < 0.27+0.056
===========================================
b. Generate a 95 % confidence interval for the difference in proportions of patients who report pain relief.
----
sample proportion: (44/120)-(21/120 = 23/120 = 0.19
standard error: 1.96*sqrt(0.19*0.81/120) = 0.07
---
95% CI: 0.19-0.07 < P < 0.19+0.07
=======================================
Cheers,
Stan H.